# Computing Wasserstein Barycenters through Gradient Flows

### v2

@@ B.3 Neural Barycenter Solvers
Based on equation 34, the insight of [visentin2025computing] is modeling $f$ and $g$ through a conditional function $f(z,s)$, $s\in\{0,1\}$. The Wasserstein barycenter problem is a straightforward generalization of the $2$ measures case. For instance, the authors consider $S=\{1,\cdots,K\}$, which means that,
$\mathbb{B}_{\mathcal{Q}}(P)=\sum_{k=1}^{K}\lambda_{k}\mathbb{W}_{p}(P,Q_{k})^{p}=\sum_{k=1}^{K}\lambda_{k}\mathbb{W}_{p}(g_{\sharp}P_{0},Q_{k})=\sum_{k=1}^{K}\lambda_{k}\lVert g-f(z,k)\rVert_{L_{p}(P_{0})}^{p},$
@@
The main challenge in [visentin2025computing] is formulating a conditional model $f:\Omega\times S$. As the authors discuss, one possibility would be using $K=|S|$ different neural nets for modeling $f(z,\cdot)$, which is similar to what is done, for instance, in the neural net methods. The authors take a different route and use conditional normalizing flows [kobyzev2020normalizing] to model these functions.
@@ B.4 Future Directions
From the previous discussion on neural net-based methods for computing Wasserstein barycenters, we can identify a few common design choices. For instance, [korotin2021continuous, kolesov2024estimating] and [gazdieva2024robust] use $\mathcal{O}(K)$ neural nets for computing Wasserstein barycenters. This design choice results in a parametrization of the underlying barycenter that grows with the number of input measures. The only neural net-based method that differs from this idea is [visentin2025computing], which uses a conditional model $f:\Omega\times S$.
@@ C.1 Calculus in Wasserstein Spaces
Notation. For the following theoretical analysis, we use the following notation. Given a measurable map $f,g:\Omega\rightarrow\Omega$. Given a measure $P\in\mathcal{P}_{2,\text{ac}}(\Omega)$, we denote,
$\lVert f\rVert_{L_{2}(P)}=\int_{\Omega}\lVert f(x)\rVert_{2}^{2}dP(x).$
@@
Given a functional $\mathbb{F}:\mathcal{P}_{2,\text{ac}}(\Omega)(\mathbb{R}^{d})\rightarrow\mathbb{R}$, we call $\delta\mathbb{F}(P):\mathbb{R}^{d}\rightarrow\mathbb{R}$, if it exists, the unique (up to additive constants) function such that
$\dfrac{d}{d\varepsilon}\mathbb{F}(P+\varepsilon\chi)\biggr|_{\epsilon=0}=\int\delta\mathbb{F}(P)d\chi$
for every perturbation $\chi$ such that, at least for $\varepsilon\in[0,\varepsilon_{0}]$, the measure $P+\varepsilon\chi\in\mathcal{P}_{p,\text{ac}}(\mathbb{R}^{d})$. This functional is called the first variation of $\mathbb{F}$ at $P$.
Following [chewi2024statistical], given a curve of densities $\{P_{t}\}_{t\geq 0}$, we can take $\chi=\partial_{t}\mathbb{F}(P_{t})$ so that $\partial_{t}\mathbb{F}(P_{t})=\int\delta\mathbb{F}(P_{t})\partial_{t}P_{t}$. From Riemannian geometry, we can define the Wasserstein gradient to $P_{t}$ as,
$\partial_{t}\mathbb{F}(P_{t})=\langle\nabla\mkern-10.0mu\nabla\mathbb{F}(P_{t}),v_{t}\rangle_{P_{t}},$ (35)
where $v_{t}\in\mathcal{T}_{P_{t}}\mathcal{P}_{p,\text{ac}}(\mathbb{R}^{d})$. Note, here, that $\nabla\mkern-10.0mu\nabla\mathbb{F}(P_{t}):\Omega\rightarrow\Omega$ is a vector field. In comparison, the first variation $\delta\mathbb{F}$ is a scalar field. These two objects are connected by the following equation [chewi2024statistical, Proposition 5.10],
$\nabla\mkern-10.0mu\nabla\mathbb{F}(P)=\nabla\delta\mathbb{F}(P),$ (36)
Meanwhile, recall that, from dynamic OT, the WGF satisfy the continuity equation,
$\partial_{t}P_{t}+\text{div}(P_{t}v_{t})=0,$
where $\text{div}(v)=\nabla\cdot v$ is the divergence of a vector field. Therefore, taking the direction of steepest descent of the functional $\mathbb{F}$, $v_{t}=-\nabla\mkern-10.0mu\nabla\mathbb{F}(P_{t})$, leads to the following connection between $\partial_{t}\mathbb{F}(P_{t})$ and $\nabla\mkern-10.0mu\nabla\mathbb{F}(P_{t})$,
$\partial_{t}\mathbb{F}(P_{t})=-\lVert\nabla\mkern-10.0mu\nabla\mathbb{F}(P_{t})\rVert_{L_{2}(P_{t})}^{2},$ (37)
@@ C.2 Wasserstein Gradient of Functionals
$\nabla\mkern-10.0mu\nabla\mathbb{V}=\nabla V.$
@@
which means that $\delta\mathbb{G}=G^{\prime}(P)$ and $\nabla\mkern-10.0mu\nabla\mathbb{G}=\nabla G^{\prime}(P(z))$.
@@
When $U(z,z^{\prime})=h(z-z^{\prime})$ for an even $h$, $\int U(z,z^{\prime})dP(z)=\int h(z-z^{\prime})P(z)dz=h\star P$, i.e., the convolution of $h$ and $P$. Furthermore, $\nabla\mkern-10.0mu\nabla\mathbb{U}=(\nabla h)\star P$
@@
$\nabla\mkern-10.0mu\nabla\mathbb{W}_{2}(P,Q)^{2}=2(\text{Id}-T_{P\rightarrow Q}^{\star})$
@@
$\nabla\mkern-10.0mu\nabla\mathbb{B}_{\mathcal{Q}}(P)$ $=\nabla\mkern-10.0mu\nabla\biggr(\sum_{k=1}^{K}\lambda_{k}\mathbb{W}_{2}(P,Q_{k})^{2}\biggr)=2\sum_{k=1}^{K}\lambda_{k}(\text{Id}-T_{P\rightarrow Q_{k}}^{\star}).$
@@
$\mathbb{W}_{2}(\hat{P},\hat{Q})^{2}=\sum_{i=1}^{n}\sum_{j=1}^{m}\gamma_{ij}^{\star}\lVert z_{i}^{(P)}-z_{j}^{(Q)}\rVert_{2}^{2}\text{, and }\mathbb{B}_{\mathcal{Q}}(\hat{P}_{\tau})=\sum_{k=1}^{K}\lambda_{k}\sum_{i=1}^{n}\sum_{j=1}^{n_{k}}\gamma_{k,i,j}^{\star}\lVert z_{\tau,i}^{(P)}-z_{j}^{(Q_{k})}\rVert_{2}^{2}.$
We can compute the Wasserstein gradient $\nabla\mkern-10.0mu\nabla\mathbb{B}_{\mathcal{Q}}$ by computing $\nabla_{z_{\tau,i}^{(P)}}\mathbb{B}_{\mathcal{Q}}$, that is,
$\nabla_{z_{\tau,i}^{(P)}}\mathbb{B}_{\mathcal{Q}}(\hat{P}_{\tau})=2\sum_{k=1}^{K}\lambda_{k}\sum_{j=1}^{n_{k}}\gamma_{k,i,j}^{\star}(z_{\tau,i}^{(P)}-z_{j}^{(Q_{k})})=\dfrac{2}{n}z_{\tau,i}^{(P)}-2\sum_{k=1}^{K}\lambda_{k}\sum_{j=1}^{K}\gamma_{k,i,j}^{\star}z_{j}^{(Q_{k})}.$
@@
$\nabla_{z_{\tau,i}^{(P)}}\mathbb{B}_{\mathcal{Q}}(\hat{P}_{\tau})=\dfrac{2}{n}\biggr(z_{\tau,i}^{(P)}-\sum_{k=1}^{K}\lambda_{k}T_{P\rightarrow Q_{k}}^{\star}(z_{\tau,i}^{(P)})\biggr),$

### v3

@@ B.3 Neural Barycenter Solvers
Based on equation 35, the insight of visentin2025computing is modeling $f$ and $g$ through a conditional function $f(z,s)$, $s\in\{0,1\}$. The Wasserstein barycenter problem is a straightforward generalization of the $2$ measures case. For instance, the authors consider $S=\{1,\cdots,K\}$, which means that,
$\mathbb{B}(P|\mathcal{Q})=\sum_{k=1}^{K}\lambda_{k}\mathbb{W}_{2}(P,Q_{k})^{2}=\sum_{k=1}^{K}\lambda_{k}\mathbb{W}_{2}(g_{\sharp}P_{0},Q_{k})^{2}=\sum_{k=1}^{K}\lambda_{k}\lVert g-f(z,k)\rVert_{L_{p}(P_{0})}^{2},$
@@
The main challenge in visentin2025computing is formulating a conditional model $f:\Omega\times S$. As the authors discuss, one possibility would be using $K=|S|$ different neural nets for modeling $f(z,\cdot)$, which is similar to what is done, for instance, in the neural net methods. The authors take a different route and use conditional normalizing flows kobyzev2020normalizing to model these functions.
@@ B.4 Future Directions
From the previous discussion on neural net-based methods for computing Wasserstein barycenters, we can identify a few common design choices. For instance, korotin2021continuous, kolesov2024estimating and gazdieva2024robust use $\mathcal{O}(K)$ neural nets for computing Wasserstein barycenters. This design choice results in a parametrization of the underlying barycenter that grows with the number of input measures. The only neural net-based method that differs from this idea is visentin2025computing, which uses a conditional model $f:\Omega\times S$.
@@ C.1 Calculus in Wasserstein Spaces
Notation. For the following theoretical analysis, we use the following notation. Given a measurable map $f,g:\Omega\rightarrow\Omega$. Given a measure $P\in\mathcal{P}_{2}(\Omega)$, we denote,
$\lVert f\rVert_{L_{2}(P)}^{2}=\int_{\Omega}\lVert f(x)\rVert_{2}^{2}dP(x).$
@@
Given a functional $\mathbb{F}:\mathcal{P}_{2}(\Omega)(\mathbb{R}^{d})\rightarrow\mathbb{R}$, we call $\delta\mathbb{F}(P):\mathbb{R}^{d}\rightarrow\mathbb{R}$, if it exists, the unique (up to additive constants) function such that
$\dfrac{d}{d\varepsilon}\mathbb{F}(P+\varepsilon\chi)\biggr|_{\epsilon=0}=\int\delta\mathbb{F}(P)d\chi,$ (36)
for every perturbation $\chi$ such that, at least for $\varepsilon\in[0,\varepsilon_{0}]$, the measure $P+\varepsilon\chi\in\mathcal{P}_{2}(\mathbb{R}^{d})$. This functional is called the first variation of $\mathbb{F}$ at $P$.
We refer readers to [chewi2024statistical, Chapter 5] and [ambrosio2008gradient, Chapter 8] for further details on gradient flows in Wasserstein spaces. The main tool we will use here is the continuity equation,
$\partial_{t}P_{t}=-\text{div}(P_{t}v_{t}),$
where the divergence operator should be understood in the measure theoretic sense. Here, $v_{t}:\Omega\rightarrow\Omega$ is a vector field. Now, given $\mathbb{F}:\mathcal{P}_{2}(\Omega)\to\mathbb{R}$ with first variation $\delta\mathbb{F}(P):\Omega\to\mathbb{R}$, the time derivative of $\mathbb{F}$ along the curve $\{P_{t}\}_{t\geq 0}$ can be computed as,
$\partial_{t}\mathbb{F}(P_{t})\overset{(1)}{=}\int\delta\mathbb{F}(P_{t})d(\partial_{t}P_{t})\overset{(2)}{=}-\int\delta\mathbb{F}(P_{t})\text{div}(P_{t}v_{t})dz\overset{(3)}{=}\int\langle\nabla\delta\mathbb{F}(P_{t}),v_{t}\rangle dP_{t}\overset{(4)}{=}\langle\nabla\delta\mathbb{F}(P_{t}),v_{t}\rangle_{P_{t}},$
where (1) is obtained by taking $\chi=\partial_{t}P_{t}$ in equation 36; (2) is obtained by applying the continuity equation; (3) is obtained by using integration by parts; finally, (4) is obtained via the definition of inner product. This motivates the equality,
$\nabla_{W}\mathbb{F}(P)=\nabla\delta\mathbb{F}(P).$ (37)
We refer readers to [chewi2024statistical, Proposition 5.10], for a more formal derivation. This discussion characterizes the Wasserstein gradient $\nabla_{W}\mathbb{F}(P):\Omega\to\Omega$ as a vector field. Therefore, taking the direction of steepest descent of the functional $\mathbb{F}$, $v_{t}=-\nabla_{W}\mathbb{F}(P_{t})$, leads to the following connection between $\partial_{t}\mathbb{F}(P_{t})$ and $\nabla_{W}\mathbb{F}(P_{t})$,
$\partial_{t}\mathbb{F}(P_{t})=-\lVert\nabla_{W}\mathbb{F}(P_{t})\rVert_{L_{2}(P_{t})}^{2},$ (38)
@@ C.2 Wasserstein Gradient of Functionals
$\nabla_{W}\mathbb{V}=\nabla V.$
@@
which means that $\delta\mathbb{G}=G^{\prime}(P)$ and $\nabla_{W}\mathbb{G}=\nabla G^{\prime}(P(z))$.
@@
When $U(z,z^{\prime})=h(z-z^{\prime})$ for an even $h$, $\int U(z,z^{\prime})dP(z)=\int h(z-z^{\prime})P(z)dz=h\star P$, i.e., the convolution of $h$ and $P$. Furthermore, $\nabla_{W}\mathbb{U}=(\nabla h)\star P$
@@
$\nabla_{W}\mathbb{W}_{2}(P,Q)^{2}=2(\text{Id}-T_{P\rightarrow Q}^{\star}).$ (39)
@@
$\nabla_{W}\mathbb{B}(P|\mathcal{Q})$ $=\nabla_{W}\biggr(\sum_{k=1}^{K}\lambda_{k}\mathbb{W}_{2}(P,Q_{k})^{2}\biggr)=2\sum_{k=1}^{K}\lambda_{k}(\text{Id}-T_{P\rightarrow Q_{k}}^{\star}).$ (40)
@@
$\mathbb{W}_{2}(\hat{P},\hat{Q})^{2}=\sum_{i=1}^{n}\sum_{j=1}^{m}\gamma_{ij}^{\star}\lVert z_{i}^{(P)}-z_{j}^{(Q)}\rVert_{2}^{2}\text{, and }\mathbb{B}(\hat{P}_{\tau}|\mathcal{Q})=\sum_{k=1}^{K}\lambda_{k}\sum_{i=1}^{n}\sum_{j=1}^{n_{k}}\gamma_{k,i,j}^{\star}\lVert z_{\tau,i}^{(P)}-z_{j}^{(Q_{k})}\rVert_{2}^{2}.$
We can compute the Wasserstein gradient $\nabla_{W}\mathbb{B}$ by computing $\nabla_{z_{\tau,i}^{(P)}}\mathbb{B}$, that is,
$\nabla_{z_{\tau,i}^{(P)}}\mathbb{B}(\hat{P}_{\tau}|\mathcal{Q})=2\sum_{k=1}^{K}\lambda_{k}\sum_{j=1}^{n_{k}}\gamma_{k,i,j}^{\star}(z_{\tau,i}^{(P)}-z_{j}^{(Q_{k})})=\dfrac{2}{n}z_{\tau,i}^{(P)}-2\sum_{k=1}^{K}\lambda_{k}\sum_{j=1}^{K}\gamma_{k,i,j}^{\star}z_{j}^{(Q_{k})}.$
@@
$\nabla_{z_{\tau,i}^{(P)}}\mathbb{B}(\hat{P}_{\tau}|\mathcal{Q})=\dfrac{2}{n}\biggr(z_{\tau,i}^{(P)}-\sum_{k=1}^{K}\lambda_{k}T_{P\rightarrow Q_{k}}^{\star}(z_{\tau,i}^{(P)})\biggr),$